OpenStudy (anonymous):

What is the sum of the interior angles of an octagon? a. 720° b. 900° c. 1080° d. 1260°

OpenStudy (anonymous):

Do you know the formula for finding the interior angles of a shape ?

OpenStudy (dan815):



OpenStudy (dan815):

look how many separate triangles you can make in an octagon

OpenStudy (dan815):

you can make 6 so 6*180 = angles in the octagon

OpenStudy (dan815):

you can always make 2 triangles less than the number of total points for example 3 points 1 triangle 4 points - square has 2 triangles = 360 degrees 5 points - pentagon has 3 triangles = 3*180=540 degrees 6 points-hexdagon has 4 triangles = 4*180 = 720 and so on

OpenStudy (dan815):

8 points, octagon - 6 triangles = 6*180 = 600 + 80 * 6 = 600+480 = 1080

OpenStudy (dan815):

make sense?

OpenStudy (anonymous):

Method 1 sum of interior angles of a polygon = (n – 2) × 180 in octagon number of sides n = 8 sum of interior angles = (8 – 2) × 180 = 1080 Method 2 sum of exterior angles of polygon is always 360° and each exterior angle = 360/n = 360/8 = 45° so so each interior angle = 180 – 45 = 135° so sum of 8 interior angles = 8 × 135 = 1080°

OpenStudy (anonymous):

Thank you!